# Conrol-Translated Finsler-type structure and Anisotropic Ginzbor–Landau models

### v1

@@ 4 Variational Control and Vortex
The first equation is a nonlinear elliptic equation for the complex field $\psi$ involving the control translation, while the second corresponds to the Ampère-type law balancing the induced current $j_{A}$ with the magnetic curvature $dA$. Both equations are defined with respect to the anisotropic divergence operator associated with the Finsler volume $d\mu$.
@@
Assume that the Finsler metric $F$ and control field $u$ satisfy $F,u\in C^{3,\alpha}$ and the Tonelli convexity conditions. Then for each fixed $\varepsilon>0$ the Euler–Lagrange system (27) admits a weak solution $(\psi,A)\in H^{1}(M;\mathbb{C})\times H^{1}(M;T^{*}M)$ minimizing $\mathcal{E}_{u}$, and the solution is smooth outside the vortex set $Z=\{x\in M:\psi(x)=0\}$.
@@
We differentiate the Euler–Lagrange equations (27) with respect to $u$. The linearized system reads
@@
where $T(a_{i})$ is the tangent vector of the unperturbed Finsler–Ginzburg–Landau vortex flow and $\mathcal{W}_{u}$ is the renormalized interaction energy from Section 3.
@@
(Conservation and Hamiltonian formulation.) In the reversible case where $u$ is divergence-free and time-independent, the vortex dynamics (30) can be written as a Hamiltonian system
@@
5 Optimal Control of the Tonelli–Finsler Energy
In this section we formulate and analyse the optimal control problem associated with the Tonelli–Finsler Ginzburg–Landau model. The control variable $u(x)$ acts as a distributed field that modifies the local anisotropy of the Lagrangian density and influences the geometry of vortices. The goal is to determine an admissible field $u^{*}$ minimizing the total energy functional
@@ 5 Optimal Control of the Tonelli–Finsler Energy
At an optimal control $u^{*}$, the stationarity condition $\delta\mathcal{J}(u^{*})[v]=0$ for all admissible $v$ yields the Euler–Lagrange system for the optimal control.
@@
(Asymptotic limit of optimal controls.) Let $u_{\varepsilon}^{*}$ denote the optimal controls minimizing $\mathcal{E}_{u}$ for fixed $\varepsilon>0$. Then, up to a subsequence,
@@
By the $\Gamma$–convergence result of Section 3, $\mathcal{E}_{u}$ converges to $\mathcal{W}_{u}$ uniformly on compact subsets of $C^{1}$ controls. Thus any cluster point $u^{*}$ of minimizers $u_{\varepsilon}^{*}$ minimizes $\mathcal{W}_{u}$. The variational inequality (39) follows by passage to the limit in (37) using weak convergence of adjoint variables and continuity of $\mathcal{W}_{u}$ in $u$ (Theorem 3.3). The convergence of energies (38) results from the $\Gamma$–liminf inequality combined with the recovery sequence property of minimizers. ∎
@@ Conclusion
This work establishes a rigorous analytical and geometric framework for Ginzburg–Landau systems within a control–translated Tonelli–Finsler setting. By embedding a smooth control field $u(x)$ as a translation in the tangent bundle, we introduced a deformed Tonelli Lagrangian that preserves Legendre duality and generates a uniformly elliptic operator $\Delta_{u}$. The corresponding energy functional $E_{u}$ remains convex, coercive, and geometrically well defined.
A detailed $\Gamma$–convergence analysis yields the renormalized energy $W_{u}$ governing vortex interactions under control translation, clarifying how the control field modifies both the Green kernel and the self-interaction potential. The results demonstrate that control acts intrinsically through geometric deformation of the underlying metric, rather than as an external force, thereby influencing both the microscopic vortex configuration and the macroscopic dynamics of the system.
The framework unifies Finsler geometry, convex analysis, and variational control into a coherent theoretical structure. It extends previous Finsler–Ginzburg–Landau models by providing a precise geometric mechanism for feedback, stabilization, and manipulation of vortices in anisotropic superconducting media. Future research directions include time-dependent Tonelli flows, optimal feedback control, and computational realizations in engineered anisotropic superconductors.

### v2

@@ 4 Variational Control and Vortex Dynamics in Tonelli-Finsler Geometry
The first equation is a nonlinear elliptic equation for the complex field $\psi$ involving the control translation, while the second corresponds to the Ampére-type law balancing the induced current $j_{A}$ with the magnetic curvature $dA$. Both equations are defined with respect to the anisotropic divergence operator associated with the Finsler volume $d\mu$.
@@
Assume that the Finsler metric $F$ and control field $u$ satisfy $F,u\in C^{3,\alpha}$ and the Tonelli convexity conditions. Then for each fixed $\varepsilon>0$ the Euler-Lagrange system (27) admits a weak solution $(\psi,A)\in H^{1}(M;\mathbb{C})\times H^{1}(M;T^{*}M)$ minimizing $\mathcal{E}_{u}$, and the solution is smooth outside the vortex set $Z=\{x\in M:\psi(x)=0\}$.
@@
We differentiate the Euler-Lagrange equations (27) with respect to $u$. The linearized system reads
@@
where $T(a_{i})$ is the tangent vector of the unperturbed Finsler-Ginzburg-Landau vortex flow and $\mathcal{W}_{u}$ is the renormalized interaction energy from Section 3.
@@
In the reversible case where $u$ is divergence-free and time-independent, the vortex dynamics (30) can be written as a Hamiltonian system
@@
5 Optimal Control of the Tonelli-Finsler Energy
In this section we formulate and analyse the optimal control problem associated with the Tonelli-Finsler-Ginzburg-Landau model. The control variable $u(x)$ acts as a distributed field that modifies the local anisotropy of the Lagrangian density and influences the geometry of vortices. The goal is to determine an admissible field $u^{*}$ minimizing the total energy functional
@@ 5 Optimal Control of the Tonelli-Finsler Energy
At an optimal control $u^{*}$, the stationarity condition $\delta\mathcal{J}(u^{*})[v]=0$ for all admissible $v$ yields the Euler-Lagrange system for the optimal control.
@@
Let $u_{\varepsilon}^{*}$ denote the optimal controls minimizing $\mathcal{E}_{u}$ for fixed $\varepsilon>0$. Then, up to a subsequence,
@@
By the $\Gamma$-convergence result of Section 3, $\mathcal{E}_{u}$ converges to $\mathcal{W}_{u}$ uniformly on compact subsets of $C^{1}$ controls. Thus any cluster point $u^{*}$ of minimizers $u_{\varepsilon}^{*}$ minimizes $\mathcal{W}_{u}$. The variational inequality (39) follows by passage to the limit in (37) using weak convergence of adjoint variables and continuity of $\mathcal{W}_{u}$ in $u$ (Theorem 3.3). The convergence of energies (38) results from the $\Gamma$-liminf inequality combined with the recovery sequence property of minimizers. ∎
@@ Conclusion
This work establishes a rigorous analytical and geometric framework for Ginzburg-Landau systems within a control-translated Tonelli-Finsler setting. By embedding a smooth control field $u(x)$ as a translation in the tangent bundle, we introduced a deformed Tonelli Lagrangian that preserves Legendre duality and generates a uniformly elliptic operator $\Delta_{u}$. The corresponding energy functional $E_{u}$ remains convex, coercive, and geometrically well defined.
A detailed $\Gamma$-convergence analysis yields the renormalized energy $W_{u}$ governing vortex interactions under control translation, clarifying how the control field modifies both the Green kernel and the self-interaction potential. The results demonstrate that control acts intrinsically through geometric deformation of the underlying metric, rather than as an external force, thereby influencing both the microscopic vortex configuration and the macroscopic dynamics of the system.
The framework unifies Finsler geometry, convex analysis, and variational control into a coherent theoretical structure. It extends previous Finsler-Ginzburg-Landau models by providing a precise geometric mechanism for feedback, stabilization, and manipulation of vortices in anisotropic superconducting media. Future research directions include time-dependent Tonelli flows, optimal feedback control, and computational realizations in engineered anisotropic superconductors.